Grammars and Lexicons: 11-721

Fall Term, 2003 Extra Credit Homework

1 Due Date

You can hand in extra credit homework any time before November 1.

2 Grading

We have not yet decided how the extra credit homework will affect your grade.

3 The Assignment: Linguistic Olympics

• The linguistic olympics is a contest for high school students. It consists of linguistic puzzles in a variety of languages. Supposedly, you do not need to know any linguistics to solve the problems, but see my notes below.

• The Linguistic Olympics web site seems to be under construction, but you can find the problems at this web page: http://www.uoregon.edu/ tpayne/lingolym/

• The problems are classified as easy, intermediate, and hard, but unfortunately you can’t see the rating because the web page is under construction.

• For extra credit, follow these steps:

– Choose any problem(s).

– Work on it as long as you want to.

– There should be a link at the bottom of each problem page that says. “Email your solution to Tom Payne and I’ll tell you how you did!” Please cc me (lsl@cs.cmu.edu) on your email and ask Tom to cc me on the reply.

– I don’t know Tom Payne, but he is probably a busy professor at University of Oregon, so please be considerate about the time it will take him to look at your answers. Please send only serious attempts.

– DO NOT LOOK AT THE ANSWER BEFORE YOU ATTEMPT THE PROBLEM.

4 Can you really do these problems without knowing linguistics?

As a sanity check, I asked some LTI faculty who are not linguists to try some of the Linguistic Olympic problems. One problem was solved quickly by everyone (five non-linguists and me). Everyone got the correct answer, but the nonlinguists used different methods than a linguist would use. Another problem was attempted by three LTI faculty — me and two non-linguists. The two non-linguists gave up before finding the solution. It was not hard for me to solve.